8. Samsung Electronics Ltd. manufactures and sells LCD monitors. The current price of a monitor is Rs. 5,000 and quantity sold is 1,000 units per year. The company believes that the price elasticity of demand is –1.5. The company decides to increase the price to Rs. 6,000. (a) How many monitors the company will be able to sell at new price? (b) What will be the effect on total revenue due to the increase in price? (c) Explain whether the increase in price is desirable or not from the viewpoint of its effect on sales and total revenue?


1
Expert's answer
2022-02-28T09:39:03-0500

a) "PED= \frac{\Delta Q}{\Delta P} \times \frac{P_1}{Q_1}"

"=\frac{Q_2-Q_1}{P_2-P_1} \times \frac{P_1}{Q_1}"

"=\frac{Q_2- 1000}{6000-5000} \times \frac{5000}{1000}"


"=\frac{Q_2- 1000}{1000} \times \frac{5000}{1000}"

"=\frac{5000Q_2- 5,000,000}{1000000} =-1.5"

"5000Q_2=-1500,000+5000000"

"5000Q_2=3500000"

"Q_2=700"

b) Revenue after price increment

"= Selling Price \times quantity sold"

"= 700\times 6000= 4,200,000"

c) Revenue before

"= 1000\times 5000= 5,000,000"

"\therefore" The increase in price is not desirable as it reduces revenue by 800,0000